# SimPro: A Simple Probabilistic Framework Towards Realistic Long-Tailed Semi-Supervised Learning

### v1

@@ 3.2 SimPro Framework
[…]
@@
  w/ ${}^{\dagger}$ (Wei & Gan, 2023) 70.9±plus-or-minus\,$\pm$0.37 76.1±plus-or-minus\,$\pm$0.42 91.9±plus-or-minus\,$\pm$0.02 92.5±plus-or-minus\,$\pm$0.19 83.2±plus-or-minus\,$\pm$0.39 85.2±plus-or-minus\,$\pm$0.12 73.8±plus-or-minus\,$\pm$0.83 79.3±plus-or-minus\,$\pm$0.30 77.6±plus-or-minus\,$\pm$0.20 79.3±plus-or-minus\,$\pm$0.48
@@
  w/ ${}^{\dagger}$ (Wei & Gan, 2023) 40.7±plus-or-minus\,$\pm$0.57 50.2±plus-or-minus\,$\pm$0.82 44.1±plus-or-minus\,$\pm$0.14 42.4±plus-or-minus\,$\pm$0.47 41.1±plus-or-minus\,$\pm$0.09 83.0±plus-or-minus\,$\pm$0.32 81.5±plus-or-minus\,$\pm$0.25
@@ 4.1 Results
We first conduct experiments on the four representative benchmark datasets with different class imbalance ratios. We denote the class imbalance ratio of labeled, unlabeled, and test data as $\gamma_{l}$, $\gamma_{u}$, and $\gamma_{t}$, respectively. Our method is compared with five competitive baseline approaches, i.e., FixMatch (Sohn et al., 2020), CReST+ (Wei et al., 2021), DASO (Oh et al., 2022), ACR (Wei & Gan, 2023), and CPE (Ma et al., 2024). Note that for a fair comparison, we first compare with ACR in the ReaLTSSL setting, where the unlabeled class distribution is unknown and inaccessible. Specifically, we compare our vanilla SimPro with ACR’s variant that removes its pre-defined anchor distributions, denoted as ${}^{\dagger}$. Then we implement ${}^{\star}$ by also alleviating the anchor distributions in our SimPro framework, comparing ${}^{\star}$ with the original ACR and CPE.
@@
${}^{\star}$ 80.0 94.1 85.0
The results of ${}^{\star}$ using anchor distributions.
To investigate the impact of the anchor distributions in ACR and CPE (Ma et al., 2024), we also incorporate them into our approach, referred to as ${}^{\star}$. Instead of calculating the distribution distance and adjusting the consistency regularization as in ACR or employing multiple anchor experts as in CPE, our usage of these anchors is more straightforward: after training for five epochs, we calculate the distance between our estimated distribution $P(y|\bm{\pi})$ and the three anchor distributions (i.e. consistent, uniform, and reversed). This calculation helps us predict the actual distribution and construct the Bayes classifier. Then we fix the marginal distribution parameters $\bm{\pi}$ in the remainder of the training.
The results in Tab. 4 indicate that (1) the usage of anchor distributions can significantly enhance the performance of ${}^{\star}$, consistently outperforming the original ACR and CPE; (2) our estimation for $\bm{\pi}$ is accurate (Fig. 5 further validates the accurate estimation of $\bm{\pi}$); (3) when the pre-defined anchors fail to cover the evaluated distributions (middle and head-tail), SimPro outperforms ACR by a large margin; (4) even compared to the original ACR, SimPro exhibits enhanced performance across all scenarios except the consistent distribution. This demonstrates the superior robustness and generalization ability of our method. We believe these advantages guarantee a better adaptable nature for SimPro in real applications and are more valuable than the accuracy improvements when using the anchor distributions.

### v2

@@ 3.2 SimPro Framework
$\text{or}\quad P(y|x;\bm{\theta})$ $=\frac{\exp(f_{\bm{\theta}}(x,y))}{\sum_{y^{\prime}}\exp(f_{\bm{\theta}}(x,y^{\prime}))}.$ (11)
@@
w/ ACR (Wei & Gan, 2023) 70.9±plus-or-minus\,$\pm$0.37 76.1±plus-or-minus\,$\pm$0.42 91.9±plus-or-minus\,$\pm$0.02 92.5±plus-or-minus\,$\pm$0.19 83.2±plus-or-minus\,$\pm$0.39 85.2±plus-or-minus\,$\pm$0.12 73.8±plus-or-minus\,$\pm$0.83 79.3±plus-or-minus\,$\pm$0.30 77.6±plus-or-minus\,$\pm$0.20 79.3±plus-or-minus\,$\pm$0.48
@@
w/ ACR (Wei & Gan, 2023) 40.7±plus-or-minus\,$\pm$0.57 50.2±plus-or-minus\,$\pm$0.82 44.1±plus-or-minus\,$\pm$0.14 42.4±plus-or-minus\,$\pm$0.47 41.1±plus-or-minus\,$\pm$0.09 83.0±plus-or-minus\,$\pm$0.32 81.5±plus-or-minus\,$\pm$0.25
@@ 4.1 Results
We first conduct experiments on the four representative benchmark datasets with different class imbalance ratios. We denote the class imbalance ratio of labeled, unlabeled, and test data as $\gamma_{l}$, $\gamma_{u}$, and $\gamma_{t}$, respectively. Our method is compared with five competitive baseline approaches, i.e., FixMatch (Sohn et al., 2020), CReST+ (Wei et al., 2021), DASO (Oh et al., 2022), ACR (Wei & Gan, 2023), and CPE (Ma et al., 2024). Note that for a fair comparison, we first compare with ACR in the ReaLTSSL setting, where the unlabeled class distribution is unknown and inaccessible. Specifically, we compare our vanilla SimPro with ACR’s variant that removes its pre-defined anchor distributions, denoted as ACR. Then we implement SimPro by also alleviating the anchor distributions in our SimPro framework, comparing SimPro with the original ACR and CPE.
@@
SimPro 80.0 94.1 85.0
The results of SimPro using anchor distributions.
To investigate the impact of the anchor distributions in ACR and CPE (Ma et al., 2024), we also incorporate them into our approach, referred to as SimPro. Instead of calculating the distribution distance and adjusting the consistency regularization as in ACR or employing multiple anchor experts as in CPE, our usage of these anchors is more straightforward: after training for five epochs, we calculate the distance between our estimated distribution $P(y|\bm{\pi})$ and the three anchor distributions (i.e. consistent, uniform, and reversed). This calculation helps us predict the actual distribution and construct the Bayes classifier. Then we fix the marginal distribution parameters $\bm{\pi}$ in the remainder of the training.
The results in Tab. 4 indicate that (1) the usage of anchor distributions can significantly enhance the performance of SimPro, consistently outperforming the original ACR and CPE; (2) our estimation for $\bm{\pi}$ is accurate (Fig. 5 further validates the accurate estimation of $\bm{\pi}$); (3) when the pre-defined anchors fail to cover the evaluated distributions (middle and head-tail), SimPro outperforms ACR by a large margin; (4) even compared to the original ACR, SimPro exhibits enhanced performance across all scenarios except the consistent distribution. This demonstrates the superior robustness and generalization ability of our method. We believe these advantages guarantee a better adaptable nature for SimPro in real applications and are more valuable than the accuracy improvements when using the anchor distributions.